# Entanglement and particle production from cosmological perturbations: a quantum optical simulation approach

### v2

@@ 2.2.1 Hamiltonian
Firstly, the Hamiltonian is constructed to track how the fields evolve. For that, we need the canonical momentum $\pi$, which is defined as the conjugate momentum to the field $v$. The canonical momentum is obtained by taking the functional derivative of the Lagrangian density with respect to the time derivative of the field, $v^{\prime}$. Up to a total derivative term, action in 31 is equivalent to albrecht1994inflation :
@@ 2.2.2 Evolution of field modes
where the quantities $r_{k}$, $\theta_{k}$, and $\phi_{k}$ are called the squeezing parameter, rotation angle, and squeezing angle, respectively, which are the functions of the conformal time. The squeezing parameter $r_{k}$ determines how much squeezing occurs. It is a quantitative measure of the amount of squeezing happening to the variables with a changing universe. The squeezing angle $\phi_{k}$ determines which variable undergoes squeezing and which undergoes stretching albrecht1994inflation . Moreover, the squeezing angle $\phi_{k}$ determines how the uncertainty is redistributed between different quadratures of a quantum state and sets the direction of squeezing in the phase space. The rotation operator is only to change the phase, which is of no consequence to our study, and we will drop it from here bhattacharyya2020cosmological . For inflation, these parameters are given as albrecht1994inflation
@@ 2.2.3 Squeezing and rotation evolution operators
Since the rotation operator only produces a global phase, we will ignore it forward. The two-mode squeezing operator applied on the vacuum state produces a squeezed vacuum state: albrecht1994inflation
@@ 2.3 Squeezing parameter for expanding and contracting backgrounds
For simplicity, we will change the variable from the conformal time to physical time using $d\eta=\frac{dt}{a(t)}$. Then, the differential equations for the squeezing parameters $r_{k}$ and $\phi_{k}$ is given by bhattacharyya2020rise
@@ 2.3.1 Expanding background
The expression for the scale factor $a(\eta)$ for the expanding background is given by bhattacharyya2020rise :
@@ 2.3.2 Contracting background
The expression for the scale factor $a(\eta)$ for the contracting background is given by bhattacharyya2020rise :
@@ 3 Continuous variable quantum computing
The concept of continuous variable quantum computing (CVQC) is based on infinite-dimensional quantum systems called qumodes. A quantum mode or qumode is an elementary unit of quantum information processing in the CVQC model whose Hilbert space corresponds to that of a quantum harmonic oscillator. The unitary operators acting on the infinite-dimensional Hilbert space are equivalent to the logic gates in this model CV_gate_synthesis ; Braunstein_2005 . This concept of using quadrature operators to form logic gates has been discussed by Lloyd and Braunstein QC_over_CV_lloyd_braunstein . The notion of universality in CVQC can be defined for transformations corresponding to Hamiltonians that are polynomial in the operators of continuous variables. A set of quantum gates is referred to as universal if any arbitrary unitary transformation can be expressed as a finite product of these gates to an arbitrary accurate approximation. So, the techniques to decompose any given unitary in terms of a universal gate set are necessary to program a quantum computer. Such methods for CVQC have been discussed in gate_decompositions . In the case of modes of the electromagnetic field, linear operations such as translations, phase shifts, squeezers, and beam splitters, in addition to a nonlinear operation such as a Kerr nonlinearity, make it possible to perform arbitrary polynomial transformations on those modes Braunstein_2005 .
A CV system ferraro2005gaussianstatescontinuousvariable ; serafini2017quantum_primer ; brask2022gaussianstates_quick_reference ; Weedbrook_2012 of $n$ modes is described by n bosons with corresponding creation and annihilation operators $\{\hat{c}^{\dagger}_{j},\hat{c}_{j}\}$ with j=1,2,,nj=1,2,...,n. These operators satisfy the canonical commutation relations, $\hat{c}_{j},\hat{c}_{k}]=[\hat{c}_{j}^{\dagger},\hat{c}_{k}^{\dagger}]=0$,and $\hat{c}_{j},\hat{c}_{k}^{\dagger}]=\delta_{jk}$, where $\delta_{jk}$ is the Kronecker delta. We now define the dimensionless quadrature operators ($\hbar$ = $1$),
@@
The continuous variable approach provides an equivalent reformulation of the density operator in terms of a quasi-probability distribution referred to as the Wigner function, defined on the phase space Weedbrook_2012 . An arbitrary quantum state $\hat{\rho}$ of an n-mode bosonic system is equivalent to a Wigner function W(r) defined over a $2n$-dimensional phase space $(\mathbb{R}^{2n},\Omega)$. The characteristic function of a quantum state $\hat{\rho}$ is defined as,
@@ 3.1 CV Gates
The unitary transformations in the Gaussian formalism must preserve the Gaussian nature of the given quantum state. This is true for the Hamiltonian, which is quadratic in the creation and annihilation operators. We consider a general quadratic Hamiltonian with complex coefficients Gaussian_states_and_beyond ,
@@
In the CV framework, we can work with Gaussian and non-Gaussian CV gates. In this paper, we only use Gaussian gates as the Hamiltonian we model is quadratic in nature, but the higher-order dynamics can also be studied using the non-Gaussian gates. The combination of Gaussian and non-Gaussian gates can be used to form a universal gate set that allows us to simulate any arbitrary Hamiltonian to desired precision. Now we present the symplectic transformations which correspond to some frequently used Gaussian unitary operations. We also describe the Kerr gate, which is a non-Gaussian gate, to provide an example of a universal gate set gate_decompositions ; Weedbrook_2012 ; brask2022gaussianstates_quick_reference ; serafini2017quantum_primer .
@@
A two-mode squeezed state can also be generated by using two single-mode squeezed states mixed by a symmetric beam splitter TMS_SMS_BS . Consider a beam splitter with transmissivity, $\delta=0.5$ then
@@
Since the action of the beam splitter is reversible, we can generate a two-mode squeezed state by combining two single-mode squeezed states using a beam splitter ferraro2005gaussianstatescontinuousvariable ; Braunstein_2005 .
@@ 3.2 Entanglement measures for gaussian states
As we have already discussed, a Gaussian state is completely characterized by its first and second moments. Consequently, the entanglement measures in the Gaussian formalism can be obtained from the covariance matrix of the given quantum state. In this section, we describe various entanglement measures such as von Neumann entropy and logarithmic negativity, along with purity and Rényi entropy for the Gaussian states. First, we briefly discuss some important concepts such as symplectic eigenvalues and partial transposition, which are essential for the calculation of the entanglement measures discussed below. For more details, refer to plenio2007_entanglement_measures ; serafini2017quantum_primer ; Plenio_entanglement ; Vidal_Werner_entanglement .
@@
Here, the set of diagonal entries is referred to as the symplectic spectrum of the covariance matrix $V$. The symplectic eigenvalues are the absolute eigenvalues of the matrix $V\Omega$. This matrix gives $2n$ eigenvalues which are purely imaginary and come in pairs $\pm iv_{i}$ where $v_{i}$ are real with i=1,2,,ni=1,2,...,n. Here we note that the symplectic eigenvalues of a covariance matrix have two-fold degeneracy. The symplectic spectrum of a covariance matrix encodes the information about the entanglement of the quantum state Brady_Agullo_Kranas_2022 ; kranas2023entanglement .
@@
The spectrum of the partially transposed matrix in Eq. 110 is independent of the local orthonormal basis and also the choice of subsystem over which the partial transposition is performed plenio2007_entanglement_measures . The covariance matrix of the partially transposed density matrix is obtained as,
@@
where $T=\mathbb{I_{2n_{1}}}\bigoplus\sigma$, with $\sigma=\bigoplus_{i=1}^{n_{2}}\sigma_{z}$, where $\sigma_{z}$ is the z-Pauli matrix. For a two-mode Gaussian state, the matrix $V^{pt}$ is obtained from the covariance matrix, $V$, by reversing the sign of all components involving momenta $p_{i}$ of the subsystem. This amounts to changing the sign of the fourth row and fourth column of the covariance matrix Brady_Agullo_Kranas_2022 .
@@ 3.2.1 Purity
In the context of Gaussian systems—especially in the presence of decoherence or thermal noise—tracking purity and related entropic measures offer a practical and informative way to assess coherence and entanglement degradation serafini2017quantum_primer .
@@
The von Neumann entropy is a reliable measure of entanglement only if the total quantum state is pure kranas2023entanglement ; Brady_Agullo_Kranas_2022 . To illustrate this, consider a two-mode Gaussian state. If the total state is pure and the subsystems are mixed, then von Neumann entropy quantifies the entanglement entropy of the state. But if the total state is mixed, the von Neumann entropy is no longer a reliable entanglement measure, as the system may or may not be entangled.
The generalization of von Neumann entropy is the quantum Rényi entropy defined as Renyi_entropy.. ,
@@ 3.2.3 Logarithmic negativity
The logarithmic negativity is easy to compute, and on Gaussian states it has an interpretation as a special type of entanglement cost. It is a reliable entanglement measure even if the total quantum state is mixed. It is important to note that $LN>0$ indicates the existence of entanglement, but $LN=0$ does not ensure separability, i.e., in general, $LN>0$ is a sufficient but not a necessary condition for entanglement. But LN does not reduce to the entropy of entanglement on pure states Plenio_entanglement ; plenio2007_entanglement_measures .
Logarithmic negativity can be computed for a bipartite Gaussian state described with a covariance matrix $V$ as serafini2017quantum_primer ; kranas2023entanglement ,
@@
where $v_{j}^{pt}$ are the symplectic eigenvalues of the covariance matrix, $V^{pt}$, of the partially transposed density matrix $\rho^{pt}$. The method to obtaine the covariance matrix $V^{pt}$ has been described above. For a two-mode system, there is only one symplectic eigenvalue that can be less than one, and we have serafini2017quantum_primer ,
@@
Consequently, for the two-mode Gaussian state, $LN>0$ is both a necessary and a sufficient condition for entanglement kranas2023entanglement .
@@ 4.1 Entanglement entropy
In order to tackle these numerical instabilities, we will rely on Rényi Entropy, which, for suitable Rényi parameters, can be more stable than von Neumann entropy. Since $S_{1}$ corresponds to the usual von Neumann entropy, finding stable Rényi parameters can be used to bound von Neumann entropy, as for any two Rényi entropies $S_{n}$ and $S_{m}$, $S_{n}\geq S_{m}$ for $n<m$. The Rényi entropy associated with cosmological perturbations is given by seroje2015effective ,
@@ 4.3 Symplectic formalism
The above analysis shows the application of the Gaussian formalism and symplectic circuit to describe the dynamics of cosmological perturbations corresponding to a quadratic Hamiltonian. This approach is also a powerful tool to include non-linearities and apply non-Gaussian states to study such interactions. This can be done by the use of Gaussian and non-Gaussian CV gates. The circuit provides a comprehensive visual representation of the dynamics of phenomena like cosmological perturbations, which can then be simulated using a Python library like QuGIT QuGIT . This allows us to study the dynamics in an efficient way, which would be very complex to perform analytically. The symplectic circuit helps us visualize the symplectic matrices applied to construct the working of the Hamiltonian. Hence, the use of techniques from continuous variable quantum optics provides a powerful framework to study analog gravity models.
@@ 5.1 Entanglement entropy
We consider an expanding cosmological background with a fixed equation of state of w=1w=-1 for the accelerating case and $w=1/3$ for the decelerating case. For an expanding, accelerating background, modes are inside the horizon at sufficiently early times $a\ll 1$, and they cross the horizon at $k\approx a$ as shown in Eq. 65. We take two values of k, $k=10^{-4}$ and $k=10^{4}$, and plot the analytic upper and lower Rényi-entropy bounds, $S_{1/2}$ and $S_{2}$, respectively, against $\ln a$, while the von Neumann entropy $S_{1}$ is obtained via a quantum optics simulation using QuGIT QuGIT .
@@
Now we consider a contracting cosmological background. For contracting acceleration, the modes begin outside the horizon for early times $a\gg 1$. Given that a contracting universe is simply the time-reverse of an expanding one, the entanglement measures in a contracting accelerating background must mirror those in an expanding decelerating background, and hence their plots coincide. In particular, at early time $a\gg 1$, both the analytic Rényi entropy and the simulated von Neumann entropy are sufficiently large, and they freeze in when the modes re-enter the horizon, as they did for expanding decelerating solutions. To avoid the numerical instability of von Neumann entropy at low $k$, we chose one of our values of $k$, a sufficiently large value, $k=10^{6}$, that eliminates the instability. With this choice, the simulated von Neumann entropy in Figure 21 produces a smooth curve that remains bounded by the corresponding Rényi entropies. The sub-graph magnifies the small region where the entropy measures for $k=10^{14}$ freeze in, as they enter the horizon, and have small oscillations before settling into their plateau.

### v3

@@ 2.2.1 Hamiltonian
Firstly, the Hamiltonian is constructed to track how the fields evolve. For that, we need the canonical momentum $\pi$, which is defined as the conjugate momentum to the field $v$. The canonical momentum is obtained by taking the functional derivative of the Lagrangian density with respect to the time derivative of the field, $v^{\prime}$. Up to a total derivative term, action in 31 is equivalent to Albrecht:1992kf :
@@ 2.2.2 Evolution of field modes
where the quantities $r_{k}$, $\theta_{k}$, and $\phi_{k}$ are called the squeezing parameter, rotation angle, and squeezing angle, respectively, which are the functions of the conformal time. The squeezing parameter $r_{k}$ determines how much squeezing occurs. It is a quantitative measure of the amount of squeezing happening to the variables with a changing universe. The squeezing angle $\phi_{k}$ determines which variable undergoes squeezing and which undergoes stretching Albrecht:1992kf . Moreover, the squeezing angle $\phi_{k}$ determines how the uncertainty is redistributed between different quadratures of a quantum state and sets the direction of squeezing in the phase space. The rotation operator is only to change the phase, which is of no consequence to our study, and we will drop it from here Bhattacharyya:2020rpy . For inflation, these parameters are given as Albrecht:1992kf
@@ 2.2.3 Squeezing and rotation evolution operators
Since the rotation operator only produces a global phase, we will ignore it forward. The two-mode squeezing operator applied on the vacuum state produces a squeezed vacuum state: Albrecht:1992kf
@@ 2.3 Squeezing parameter for expanding and contracting backgrounds
For simplicity, we will change the variable from the conformal time to physical time using $d\eta=\frac{dt}{a(t)}$. Then, the differential equations for the squeezing parameters $r_{k}$ and $\phi_{k}$ is given by Bhattacharyya:2020kgu
@@ 2.3.1 Expanding background
The expression for the scale factor $a(\eta)$ for the expanding background is given by Bhattacharyya:2020kgu :
@@ 2.3.2 Contracting background
The expression for the scale factor $a(\eta)$ for the contracting background is given by Bhattacharyya:2020kgu :
@@ 3 Continuous variable quantum computing
The concept of continuous variable quantum computing (CVQC) is based on infinite-dimensional quantum systems called qumodes. A quantum mode or qumode is an elementary unit of quantum information processing in the CVQC model whose Hilbert space corresponds to that of a quantum harmonic oscillator. The unitary operators acting on the infinite-dimensional Hilbert space are equivalent to the logic gates in this model Cochran:2024cri ; Braunstein:2005zz . This concept of using quadrature operators to form logic gates has been discussed by Lloyd and Braunstein Lloyd:1998jk . The notion of universality in CVQC can be defined for transformations corresponding to Hamiltonians that are polynomial in the operators of continuous variables. A set of quantum gates is referred to as universal if any arbitrary unitary transformation can be expressed as a finite product of these gates to an arbitrary accurate approximation. So, the techniques to decompose any given unitary in terms of a universal gate set are necessary to program a quantum computer. Such methods for CVQC have been discussed in Kalajdzievski:2021axm . In the case of modes of the electromagnetic field, linear operations such as translations, phase shifts, squeezers, and beam splitters, in addition to a nonlinear operation such as a Kerr nonlinearity, make it possible to perform arbitrary polynomial transformations on those modes Braunstein:2005zz .
A CV system Ferraro:2005hen ; Serafini:2017rrn ; Brask:2021kvs ; Weedbrook:2011wxo of $n$ modes is described by n bosons with corresponding creation and annihilation operators $\{\hat{c}^{\dagger}_{j},\hat{c}_{j}\}$ with j=1,2,,nj=1,2,...,n. These operators satisfy the canonical commutation relations, $\hat{c}_{j},\hat{c}_{k}]=[\hat{c}_{j}^{\dagger},\hat{c}_{k}^{\dagger}]=0$,and $\hat{c}_{j},\hat{c}_{k}^{\dagger}]=\delta_{jk}$, where $\delta_{jk}$ is the Kronecker delta. We now define the dimensionless quadrature operators ($\hbar$ = $1$),
@@
The continuous variable approach provides an equivalent reformulation of the density operator in terms of a quasi-probability distribution referred to as the Wigner function, defined on the phase space Weedbrook:2011wxo . An arbitrary quantum state $\hat{\rho}$ of an n-mode bosonic system is equivalent to a Wigner function W(r) defined over a $2n$-dimensional phase space $(\mathbb{R}^{2n},\Omega)$. The characteristic function of a quantum state $\hat{\rho}$ is defined as,
@@ 3.1 CV Gates
The unitary transformations in the Gaussian formalism must preserve the Gaussian nature of the given quantum state. This is true for the Hamiltonian, which is quadratic in the creation and annihilation operators. We consider a general quadratic Hamiltonian with complex coefficients Adesso:2014npz ,
@@
In the CV framework, we can work with Gaussian and non-Gaussian CV gates. In this paper, we only use Gaussian gates as the Hamiltonian we model is quadratic in nature, but the higher-order dynamics can also be studied using the non-Gaussian gates. The combination of Gaussian and non-Gaussian gates can be used to form a universal gate set that allows us to simulate any arbitrary Hamiltonian to desired precision. Now we present the symplectic transformations which correspond to some frequently used Gaussian unitary operations. We also describe the Kerr gate, which is a non-Gaussian gate, to provide an example of a universal gate set Kalajdzievski:2021axm ; Weedbrook:2011wxo ; Brask:2021kvs ; Serafini:2017rrn .
@@
A two-mode squeezed state can also be generated by using two single-mode squeezed states mixed by a symmetric beam splitter Wolf:2002bhs . Consider a beam splitter with transmissivity, $\delta=0.5$ then
@@
Since the action of the beam splitter is reversible, we can generate a two-mode squeezed state by combining two single-mode squeezed states using a beam splitter Ferraro:2005hen ; Braunstein:2005zz .
@@ 3.2 Entanglement measures for gaussian states
As we have already discussed, a Gaussian state is completely characterized by its first and second moments. Consequently, the entanglement measures in the Gaussian formalism can be obtained from the covariance matrix of the given quantum state. In this section, we describe various entanglement measures such as von Neumann entropy and logarithmic negativity, along with purity and Rényi entropy for the Gaussian states. First, we briefly discuss some important concepts such as symplectic eigenvalues and partial transposition, which are essential for the calculation of the entanglement measures discussed below. For more details, refer to Plenio:2007zz ; Serafini:2017rrn ; Plenio:2005cwa ; Vidal:2002zz .
@@
Here, the set of diagonal entries is referred to as the symplectic spectrum of the covariance matrix $V$. The symplectic eigenvalues are the absolute eigenvalues of the matrix $V\Omega$. This matrix gives $2n$ eigenvalues which are purely imaginary and come in pairs $\pm iv_{i}$ where $v_{i}$ are real with i=1,2,,ni=1,2,...,n. Here we note that the symplectic eigenvalues of a covariance matrix have two-fold degeneracy. The symplectic spectrum of a covariance matrix encodes the information about the entanglement of the quantum state Brady:2022ffk ; Kranas:2023aph .
@@
The spectrum of the partially transposed matrix in Eq. 110 is independent of the local orthonormal basis and also the choice of subsystem over which the partial transposition is performed Plenio:2007zz . The covariance matrix of the partially transposed density matrix is obtained as,
@@
where $T=\mathbb{I_{2n_{1}}}\bigoplus\sigma$, with $\sigma=\bigoplus_{i=1}^{n_{2}}\sigma_{z}$, where $\sigma_{z}$ is the z-Pauli matrix. For a two-mode Gaussian state, the matrix $V^{pt}$ is obtained from the covariance matrix, $V$, by reversing the sign of all components involving momenta $p_{i}$ of the subsystem. This amounts to changing the sign of the fourth row and fourth column of the covariance matrix Brady:2022ffk .
@@ 3.2.1 Purity
In the context of Gaussian systems—especially in the presence of decoherence or thermal noise—tracking purity and related entropic measures offer a practical and informative way to assess coherence and entanglement degradation Serafini:2017rrn .
@@
The von Neumann entropy is a reliable measure of entanglement only if the total quantum state is pure Kranas:2023aph ; Brady:2022ffk . To illustrate this, consider a two-mode Gaussian state. If the total state is pure and the subsystems are mixed, then von Neumann entropy quantifies the entanglement entropy of the state. But if the total state is mixed, the von Neumann entropy is no longer a reliable entanglement measure, as the system may or may not be entangled.
The generalization of von Neumann entropy is the quantum Rényi entropy defined as Li:2022ktm ,
@@ 3.2.3 Logarithmic negativity
The logarithmic negativity is easy to compute, and on Gaussian states it has an interpretation as a special type of entanglement cost. It is a reliable entanglement measure even if the total quantum state is mixed. It is important to note that $LN>0$ indicates the existence of entanglement, but $LN=0$ does not ensure separability, i.e., in general, $LN>0$ is a sufficient but not a necessary condition for entanglement. But LN does not reduce to the entropy of entanglement on pure states Plenio:2005cwa ; Plenio:2007zz .
Logarithmic negativity can be computed for a bipartite Gaussian state described with a covariance matrix $V$ as Serafini:2017rrn ; Kranas:2023aph ,
@@
where $v_{j}^{pt}$ are the symplectic eigenvalues of the covariance matrix, $V^{pt}$, of the partially transposed density matrix $\rho^{pt}$. The method to obtaine the covariance matrix $V^{pt}$ has been described above. For a two-mode system, there is only one symplectic eigenvalue that can be less than one, and we have Serafini:2017rrn ,
@@
Consequently, for the two-mode Gaussian state, $LN>0$ is both a necessary and a sufficient condition for entanglement Kranas:2023aph .
@@ 4.1 Entanglement entropy
In order to tackle these numerical instabilities, we will rely on Rényi Entropy, which, for suitable Rényi parameters, can be more stable than von Neumann entropy. Since $S_{1}$ corresponds to the usual von Neumann entropy, finding stable Rényi parameters can be used to bound von Neumann entropy, as for any two Rényi entropies $S_{n}$ and $S_{m}$, $S_{n}\geq S_{m}$ for $n<m$. The Rényi entropy associated with cosmological perturbations is given by Seroje:2015tsa ,
@@ 4.3 Symplectic formalism
The above analysis shows the application of the Gaussian formalism and symplectic circuit to describe the dynamics of cosmological perturbations corresponding to a quadratic Hamiltonian. This approach is also a powerful tool to include non-linearities and apply non-Gaussian states to study such interactions. This can be done by the use of Gaussian and non-Gaussian CV gates. The circuit provides a comprehensive visual representation of the dynamics of phenomena like cosmological perturbations, which can then be simulated using a Python library like QuGIT Brandao:2022voc . This allows us to study the dynamics in an efficient way, which would be very complex to perform analytically. The symplectic circuit helps us visualize the symplectic matrices applied to construct the working of the Hamiltonian. Hence, the use of techniques from continuous variable quantum optics provides a powerful framework to study analog gravity models.
@@ 5.1 Entanglement entropy
We consider an expanding cosmological background with a fixed equation of state of w=1w=-1 for the accelerating case and $w=1/3$ for the decelerating case. For an expanding, accelerating background, modes are inside the horizon at sufficiently early times $a\ll 1$, and they cross the horizon at $k\approx a$ as shown in Eq. 65. We take two values of k, $k=10^{-4}$ and $k=10^{4}$, and plot the analytic upper and lower Rényi‑entropy bounds, $S_{1/2}$ and $S_{2}$, respectively, against $\ln a$, while the von Neumann entropy $S_{1}$ is obtained via a quantum optics simulation using QuGIT Brandao:2022voc .
@@
Now we consider a contracting cosmological background. For contracting acceleration, the modes begin outside the horizon for early times $a\gg 1$. Given that a contracting universe is simply the time‐reverse of an expanding one, the entanglement measures in a contracting accelerating background must mirror those in an expanding decelerating background, and hence their plots coincide. In particular, at early time $a\gg 1$, both the analytic Rényi entropy and the simulated von Neumann entropy are sufficiently large, and they freeze in when the modes re-enter the horizon, as they did for expanding decelerating solutions. To avoid the numerical instability of von Neumann entropy at low $k$, we chose one of our values of $k$, a sufficiently large value, $k=10^{6}$, that eliminates the instability. With this choice, the simulated von Neumann entropy in Figure 21 produces a smooth curve that remains bounded by the corresponding Rényi entropies. The sub-graph magnifies the small region where the entropy measures for $k=10^{14}$ freeze in, as they enter the horizon, and have small oscillations before settling into their plateau.